# Reducing Size Bias in Sampling Infectious Disease Spread on Networks

### v2

@@ 2.3 SIR model
where $\beta/N$ is the transmission rate per infected individual per unit time and $\gamma$ is the recovery rate per infected individual per unit time.
The result of these two reactions/events is either the, respective, increase or decrease in the number of individuals in states $\{S,I,R\}$, which is translated as the change in the state vector $\bm{X}$. Let $\bm{v}_{i}\equiv(v_{S},v_{I},v_{R})$ be the state-change vector, where $i=1$ denotes reaction (5) and $i=2$ denotes reaction (6). The state change vector for reactions (5) and (6) are $\bm{v}_{1}=(-1,1,0)$, and $\bm{v}_{2}=(0,-1,1)$, respectively.
@@
$\quad-\underbrace{a_{j}(\bm{X})P(\bm{X},t|\bm{X}_{0},t_{0})}_{\text{outflow rate, probability to move away from state }\bm{X}\text{ given that }\bm{X}(t)=\bm{X}-\bm{v}_{j}}\Bigg{]},$ (7)
@@
Due to the nonlinearity of Eq. (2.3) obtaining a closed-form solution is not possible. However, we can simulate the SIR model on networks using the Gillespie algorithm (Gillespie, 2007), which provides exact numerical realizations of the system dynamics.
2.4 Network Generation and SIR Model Simulation
Above, we have discussed the theoretical aspects of network types, network properties, sampling algorithms, and a stochastic SIR model. In this section, we provide the libraries used for generating the networks and for simulating the SIR model on networks. The pseudocode for the RW and the MHRW sampling algorithms is in Appendix A.
We use the NetworkX Python library (Hagberg et al., 2008) to generate the three networks described in Section 2.1.1. ER networks are generated using the gnp_random_graph function, SW networks with the watts_strogatz_graph function, and SF networks with the configuration_model function, where the input is a sequence of degree values drawn from a power-law distribution (using the random.zipf function from the NumPy library (Harris et al., 2020)). We generate $n=10,000$ networks for each type using the parameters listed in Table 1 to achieve the precision of $10^{-2}$ in the mean estimates of network characteristics, based on the relationship between sample size (or the number of simulations) and the standard error of the mean (Bondy and Zlot, 1976), given by
@@ 2.4 Network Generation and SIR Model Simulation
Table 1: Chosen parameters for generating examples of three (ER, SW, SF) networks using the NetworkX (Hagberg et al., 2008) library.
In Figure 2, examples of an ER network, a SW network, and an SF network are shown, each generated using the NetworkX library with $50$ nodes, an average degree of $5$, and other parameters as mentioned in Table 1.
We use the EoN Python library (Miller and Ting, 2020) to implement the Gillespie algorithm (Gillespie, 2007) simulating a stochastic SIR model (Section 2.3) on a network. The parameters for the SIR model include a recovery rate, $\gamma\in\{1,1/7,1/14\}$, and a transmission rate, $\beta/N$, which ranges from 00 to $1$ with a step size of $0.1$. We first consider $\gamma=1$, so that $\beta/N$ can be interpreted as a basic reproduction number for the SIR model. To understand the impact of longer recovery periods on estimates from samples, we use $\gamma=1/7$ and $1/14$, which correspond to average recovery periods for influenza (one week) (Chowell et al., 2008) and COVID-19 (two weeks) (SeyedAlinaghi et al., 2021), respectively.
After simulating the SIR model on the three networks, we obtain $100$ samples with $500$ nodes for each network type (ER, SW, and SF), using the RW and MHRW sampling algorithms (see pseudocode in Appendices 1,2). We generate $2500$ nodes per sample for each algorithm and discard the initial $2000$ nodes as burn-in nodes; to allow the sampling algorithms to reach their stationary distribution and reduce the bias from the starting node. Here, the sample size of $500$ (equivalent to $5\%$ of the population) is chosen to keep the margin of error approximately ±$\pm 5\%$ in the estimates (Conroy et al., 2016). We obtain $100$ samples from each of the $10,000$ networks in each network type, to get the precision of mean estimates up to the third decimal place (Bondy and Zlot, 1976). The code for the entire process pipeline is available on GitHub.
@@ 2.5 Hypothesis Test Statements
We perform hypothesis testing to statistically compare the estimates of disease metrics in RW and MHRW samples, relative to the underlying network (UN). The first step is to check if the data is normally distributed. We use one-sample Kolmogorov-Smirnov (KS) test (Jr, 1951) to test the following hypotheses:
@@
The proportion of infected nodes, for all $\beta$ values, follow a Normal distribution.
@@
The proportion of infected nodes, for all $\beta$ values, does not follow a Normal distribution.
@@
The number of secondary infections, for all $\beta$ values, follow a Normal distribution.
@@
The number of secondary infections, for all $\beta$ values, does not follow a Normal distribution.
Based on the results of the normality test, we then use the one-tailed Mann-Whitney U (U) test (Mann and Whitney, 1947) to test the hypotheses for comparing two samples and understand the direction of estimates from the samples relative to the UN. These hypotheses are stated as,
@@
    There is no difference between the probability distribution of the proportion of infected nodes in RW, MHRW and UN.
@@
    There is no difference between the distribution of the proportion of infected nodes in RW or MHRW and UN
@@
    There is no difference between the probability distribution of the number of secondary infections in RW, MHRW and UN
@@
    There is no difference between the distribution of the number of secondary infections in RW or MHRW and UN
@@ 3 Results
In this section, we present a comparative analysis of the RW and MHRW sampling methods—focussing on their ability to capture the degree distribution of the UN and their impact on the estimation of disease characteristics for three network types (ER, SW, and SF) relative to the UN.
Since RW and MHRW are memoryless algorithms, resulting in duplicate nodes in a sample (Hu and Lau, 2013). We compare the outcomes when duplicates are removed or retained. This analysis aims to highlight how this subtle aspect can significantly influence the accuracy of disease metric estimation across network types.
@@ 3.1 Degree Distribution
Figure 3: The ratio of degree distributions for samples generated by the RW algorithm (red diamonds) and the MHRW algorithm (blue triangles), relative to the UN, is presented for three network types (ER, SW, SF). The UN degree distribution is based on SIR model simulations ($\gamma=1$) for $10,000$ networks of each type (parameters in Table 1) generated with NetworkX. Sample degree distribution is derived from $100$ samples (size $500$) for $10,000$ networks using RW and MHRW.
@@
In Figure 3(c), both sampling methods do not perform well for the SF network. There is significant variability and both methods over-represent high degree nodes. The RW samples consistently show a higher ratio than the MHRW samples, showing more size bias. The presence of “hubs” in SF networks leads to a high variance in the degree distribution, leading to higher variability in the ratio compared to the ER and SW networks.
These results suggest that for networks with properties similar to the ER and SW, the MHRW sampling algorithm may be a more suitable choice for unbiased degree distribution. Whereas networks with SF properties require a more delegated approach, such as SLSR (Jiao, 2024), where core and peripheral nodes are sampled separately for balanced sampling.
@@ 3.2 Disease metrics
In the previous section, we compared the degree distribution of the samples generated using the RW and MHRW algorithms; we observed that the degree distribution of MHRW samples is closer to the UN than RW samples, due to bias towards high-degree nodes. Now, we compare the sampling algorithms in estimating three disease metrics: 1) proportion of infected nodes, 2) average number of secondary infections, and 3) time-to-infection. We simulate the SIR model on networks to estimate the disease metrics in the samples and the UN. Here, we present the results for $\gamma=1$; see C.2 for $\gamma\in\{1/7,1/14\}$.
@@ 3.2.1 Proportion of infected nodes
Figure 4: The ratio of the proportion of infected nodes during an epidemic between RW (dashed red) and MHRW (dotted blue) samples, relative to their respective UNs, is shown for (1) retained and (2) removed duplicate nodes. The light red and blue bands indicate the range of one standard deviation. The UN infection count is based on SIR model simulations ($\gamma=1$) for $10,000$ networks of each type (parameters in Table 1) generated with NetworkX (Hagberg et al., 2008). Sample infection counts are derived from $100$ samples (size $500$) for $10,000$ networks using RW and MHRW.
@@
For $\beta>0.2$, RW samples consistently overestimate the proportion of infected nodes in ER and SW networks compared to MHRW samples, regardless of the inclusion of duplicate nodes (Figure 4). In SF networks, while MHRW samples tend to overestimate the average number of infected nodes relative to RW samples, estimates from both algorithms exhibit substantial variability and fail to accurately represent the underlying network structure.
For ER networks, the MHRW samples overestimate the proportion of infected nodes when duplicate nodes are removed (Figure 3(g)) but aligns well when duplicate nodes are retained (Figure 3(d)). RW estimates remain consistent regardless of duplicate removal. Similar behaviour is observed for SW networks in Figure 3(h) and Figure 3(e).
@@
We use hypothesis testing to statistically compare the estimates from the sampling algorithms. First, we conduct a one-sample KS test to assess whether the proportion of infected nodes follows a Normal distribution, as stated in Hypothesis (2.1), with significance level $\alpha=0.05$. The KS test statistic is summarized in B Table 5 for three data groups UN, RW, and MHRW for ER, SW, and SF networks. The KS statistic is apporximately $0.5$ across networks and data groups, with all corresponding p-values less than $0.05$. These results indicate significant deviation from normality, providing evidence to reject $H_{0}$ in Hypothesis (2.1).
Since the proportion of infected nodes does not follow a Normal distribution, we conduct a one-tailed U-test to compare the RW and the MHRW sample estimates with the UN. We test two hypotheses, one for statistically significant overestimation, Hypothesis (2.3), and another for statistically significant underestimation Hypothesis (2.4), with significance level $\alpha=0.05$. The U-test statistic provides a measure of closeness between the sample estimates and the UN, with a relatively larger value indicating greater similarity to the UN. The statistics of the U-test are summarized in Table 2 for samples with duplicate nodes and samples without duplicate nodes.
The p-values are less than $0.05$ across network types for the RW samples (with and without duplicates) and the MHRW samples without duplicates nodes, for hypothesis (2.3). This provides evidence to reject $H_{0}$ (2.3) and indicates the significant overestimation of the proportion of infected nodes. For the MHRW samples with duplicate nodes, the p-value is less than $0.05$ for the SF network, evidence to reject $H_{0}$ in Hypothesis (2.3). This indicates the significant overestimation by the MHRW samples for the SF network. On the other hand, for the MHRW samples with duplicate nodes in the ER and SW networks, the p-values are less than $0.05$ for Hypothesis (2.4) (reject $H_{0}$), indicating significant underestimation of the proportion of infected nodes.
@@
Table 2: U-test statistics (Hypotheses (2.3) and (2.4)) for comparing the estimates of the proportion of infected nodes in the samples generated using the RW and the MHRW sampling algorithms with the UN, with and without duplicate nodes.
@@ 3.2.2 Average secondary infections
Figure 5: The ratio of the number of secondary infections during an epidemic between RW (dashed red) and MHRW (dotted blue) samples, relative to their respective UNs, is shown for (1) retained and (2) removed duplicate nodes. The light red and blue bands indicate the range of one standard deviation. The UN estimates are based on SIR model simulations ($\gamma=1$) for $10,000$ networks of each type (parameters in Table 1) generated with NetworkX (Hagberg et al., 2008). Sample estimates are derived from $100$ samples (size $500$) for $10,000$ networks using RW and MHRW.
@@
For $\beta>0.3$, RW samples (with, or without, duplicates) overestimate the number of secondary infections in ER and SW networks compared to the MHRW samples as shown in Figures 4(a), 4(b), 4(d), and 4(e).
@@
For ER and SW networks, the removal of duplicate nodes results in overestimation by MHRW samples (Figure 4(d)) while it aligns well when duplicate nodes are retained (Figure 4(a)). However, no change is observed for RW estimates in both scenarios (with or without duplicate nodes). Similar behaviour is observed for SW networks in Figure 4(e) and Figure 4(b).
@@
For the SF network, both sampling algorithms overestimate the number of secondary infections, with high variability (Figures 4(c) and 4(f)). However, when duplicate nodes are retained, MHRW sample estimates are closer to the UN than the RW estimates, whereas this behaviour reverses when duplicates are removed. We consistently observe that the network properties and disease metrics are highly variable and more sensitive to changes in the system parameters for the SF networks, indicating the limitation of both sampling algorithms to represent the UN well.
@@
We conduct hypothesis testing to statistically compare the estimates from the sampling algorithms. First, we conduct the one-sample KS test to assess whether the number of secondary infections follows a Normal distribution, as stated in Hypothesis (2.2), with significance level $\alpha=0.05$. The KS test statistic is summarized in B Table 5 for three data groups that are UN, RW, and MHRW for ER, SW, and SF networks. The KS statistic ranges from $0.5$ to approximately $0.6$ across networks and data groups, with all corresponding p-values less than $0.05$. These results indicate significant deviation from normality, providing evidence to reject $H_{0}$ in Hypothesis (2.2).
@@
Since the number of secondary infections does not follow a Normal distribution, we conduct a one-tailed U-test to compare the RW and the MHRW sample estimates with the UN. We test two hypotheses, one for statistically significant overestimation (Hypothesis (2.5)), and another for statistically significant underestimation (Hypothesis (2.6)), with significance level $\alpha=0.05$. The statistics of the U-test are summarized in Table 3 for samples with duplicate nodes, and samples without duplicate nodes.
@@
The p-values are less than $0.05$ across network types for the RW samples (with and without duplicates) and the MHRW samples without duplicates nodes, for hypothesis (2.5). This provides evidence to reject $H_{0}$ in Hypothesis (2.5) and indicates a significant overestimation of the number of secondary infections. For the MHRW samples with duplicate nodes, p-values are less than $0.05$ for the ER and SF networks, evidence to reject $H_{0}$ in Hypothesis (2.6), implying a significant underestimation of the number of secondary infections. However, for the SW network, the p-value is greater than $0.05$ (accept $H_{0}$) for the MHRW sample estimates with duplicate nodes, for Hypotheses (2.6) and (2.5), indicating that the MHRW estimates are statistically similar to the UN.
@@
Similar to the proportion of infected nodes, in the case of the number of secondary infections the U-test statistics (see Table 3) are higher for the MHRW sample estimates (with or without duplicates) compared to the RW for the ER and SW networks, implying that MHRW estimates are closer to the UN. For the SF networks, the RW sample estimates are closer to the UN compared to the MHRW, indicated by the higher values of U-test statistics in Table 3.

### v3

@@ 2.3 Stochastic SIR model
where $\beta/N$ is the transmission rate per infected individual per unit of time, and $\gamma$ is the recovery rate per infected individual per unit of time.
The result of these two reactions/events is either the respective increase or decrease in the number of individuals in states $\{S,I,R\}$, which is translated as the change in the state vector $\bm{X}$. Let $\bm{v}_{i}\equiv(v_{S},v_{I},v_{R})$ be the state-change vector, where $i=1$ denotes reaction (5) and $i=2$ denotes reaction (6). The state change vector for reactions (5) and (6) are $\bm{v}_{1}=(-1,1,0)$, and $\bm{v}_{2}=(0,-1,1)$, respectively.
@@
$\quad-\underbrace{a_{j}(\bm{X})P(\bm{X},t|\bm{X}_{0},t_{0})}_{\text{outflow rate, probability to move away from state }\bm{X}\text{ given that }\bm{X}(t)=\bm{X}}\Bigg{]},$ (7)
@@
Due to the nonlinearity of Eq. (2.3), obtaining a closed-form solution is not possible. However, we can simulate the SIR model on networks using the Gillespie algorithm (Gillespie, 2007), which provides exact numerical realizations of the system dynamics.
2.4 Network generation and stochastic SIR model simulations
Above, we have discussed the theoretical aspects of network types, network properties, sampling algorithms, and a stochastic SIR model. In this section, we provide the libraries used for generating the networks and for simulating the SIR model on networks. The pseudocode for the RW and the MHRW sampling algorithms is in A.
We use the NetworkX Python library (Hagberg et al., 2008) to generate the three networks described in Section 2.1.1. ER networks are generated using the gnp_random_graph function, SW networks with the watts_strogatz_graph function, and SF networks with the configuration _model function, where the input is a sequence of degree values drawn from a power-law distribution (using the random.zipf function from the NumPy library (Harris et al., 2020)). We generate $n=10,000$ networks for each type using the parameters listed in Table 1 to achieve a precision of $10^{-2}$ in the mean estimates of network characteristics, based on the relationship between sample size (or the number of simulations) and the standard error of the mean (Bondy and Zlot, 1976), given by
@@ 2.4 Network generation and stochastic SIR model simulations
Table 1: Chosen parameters for generating examples of three (ER, SW, SF) networks using the NetworkX library (Hagberg et al., 2008).
In Figure 2, examples of an ER network, a SW network, and a SF network are shown, each generated using the NetworkX library with $50$ nodes, an average degree of $5$, and other parameters as mentioned in Table 1.
We use the EoN Python library (Miller and Ting, 2020) to implement the Gillespie algorithm (Gillespie, 2007) for simulating a stochastic SIR model (Section 2.3) on a network. The parameters for the SIR model include a recovery rate, $\gamma\in\{1,1/7,1/14\}$, and a transmission rate, $\beta/N$, which ranges from 00 to $1$ with a step size of $0.1$. We first consider $\gamma=1$ so that $\beta/N$ can be interpreted as a basic reproduction number for the SIR model. To understand the impact of longer recovery periods on estimates from samples, we use $\gamma=1/7$ and $1/14$, which correspond to average recovery periods for influenza (one week) (Chowell et al., 2008) and COVID-19 (two weeks) (SeyedAlinaghi et al., 2021), respectively.
After simulating the SIR model on the three networks, we obtain $100$ samples with $500$ nodes for each network type (ER, SW, and SF), using the RW and MHRW sampling algorithms (see pseudocode in Appendices 1,2). We generate $2500$ nodes per sample for each algorithm and discard the initial $2000$ nodes as burn-in nodes to allow the sampling algorithms to reach their stationary distribution and minimize the influence of the initial node on the resulting samples. Here, the sample size of $500$ (equivalent to $5\%$ of the population) is chosen to keep the margin of error approximately ±$\pm 5\%$ in the estimates (Conroy et al., 2016). We obtain $100$ samples from each of the $10,000$ networks in each network type to get the precision of mean estimates up to the third decimal place (Bondy and Zlot, 1976). The code for the entire process pipeline is available on GitHub.
@@ 2.5 Hypothesis test statements
We perform hypothesis testing to statistically compare the estimates of disease metrics in RW and MHRW samples relative to the underlying network (UN). The first step is to check if the data is normally distributed. We use one-sample Kolmogorov-Smirnov (KS) test (Jr, 1951) to test the following hypotheses for the ER, SW, and SF networks, respectively:
@@
The proportion of infected nodes for all $\beta$ values follows a Normal distribution.
@@
The proportion of infected nodes for all $\beta$ values does not follow a Normal distribution.
@@
The number of secondary infections for all $\beta$ values follows a Normal distribution.
@@
The number of secondary infections for all $\beta$ values does not follow a Normal distribution.
@@
Time-to-infection, for all $\beta$ values, follows a Normal distribution.
Alternative Hypothesis (H1):
Time-to-infection, for all $\beta$ values, does not follow a Normal distribution.
Based on the results of the normality test, we then use the one-tailed Mann-Whitney U (U) test (Mann and Whitney, 1947) to test the hypotheses for comparing two samples and understand the direction of estimates from the samples relative to the UN. We test the following hypotheses for the ER, SW, and SF networks, respectively:
Hypothesis 2.4.
Null Hypothesis (H0):
@@
    There is no difference between the probability distribution of the proportion of infected nodes in RW and UN or MHRW and UN.
@@
    There is no difference between the distribution of the proportion of infected nodes in RW and UN or MHRW and UN
@@
    There is no difference between the probability distribution of the number of secondary infections in RW and UN or MHRW and UN
@@
    There is no difference between the distribution of the number of secondary infections in RW and UN or MHRW and UN
@@
Hypothesis 2.8.
Null Hypothesis (H0):
  • 1.
    There is no difference between the probability distribution of time-to-infection in RW and UN or MHRW and UN
  • 2.
    Time-to-infection for UN is greater than or equal to that in RW or MHRW.
Alternative Hypothesis (H1):
Time-to-infection in the UN tends to be less than that in RW or MHRW.
Hypothesis 2.9.
Null Hypothesis (H0):
  • 1.
    There is no difference between the distribution of time-to-infection in RW and UN or MHRW and UN
  • 2.
    Time-to-infection in UN is less than or equal to those in RW or MHRW.
Alternative Hypothesis (H1):
Time-to-infection in the UN tends to be greater than that in RW or MHRW.
@@ 3 Results
In this section, we present a comparative analysis of the RW and MHRW sampling algorithms, focussing on their ability to capture the degree distribution of the UN and on their impact on the estimation of disease metrics for three network types (ER, SW, and SF) relative to the UN.
Since RW and MHRW are memoryless algorithms, they yield duplicate nodes in a sample (Hu and Lau, 2013). We compare the outcomes when duplicates are removed or retained. This analysis aims to highlight how this subtle aspect can significantly influence the accuracy of disease metric estimation across network types.
@@ 3.1 Degree distribution
Figure 3: The ratio of degree distributions for samples generated by the RW algorithm (red diamonds) and the MHRW algorithm (blue triangles), relative to the UN, is presented for three network types (ER, SW, SF). The UN degree distribution is based on $10,000$ networks of each type (the parameters are found in Table 1) generated with the NetworkX library (Hagberg et al., 2008). Sample degree distribution is derived from $100$ samples (size $500$) for $10,000$ networks using RW and MHRW.
@@
In Figure 3(c), both sampling methods do not perform well for the SF network. There is significant variability, and both methods over-represent high-degree nodes. The RW samples consistently show a higher ratio than the MHRW samples, showing more size bias. The presence of “hubs” in SF networks leads to a high variance in the degree distribution, leading to higher variability in the ratio compared to the ER and SW networks.
These results suggest that for networks with properties similar to the ER and SW, the MHRW sampling algorithm may be a more suitable choice for unbiased degree distribution. At the same time, networks with SF properties require a more delegated approach, such as Sampling for Large-Scale Networks at Low Sampling Rates (SLSR) (Jiao, 2024), where core and peripheral nodes are sampled separately for balanced sampling.
@@ 3.2 Disease metrics
In the previous section, we compared the degree distributions of the samples generated using the RW and MHRW algorithms. We observed that the degree distribution of MHRW samples is closer to that of the UN than that of RW samples due to a bias towards high-degree nodes. Now, we compare the sampling algorithms in estimating three disease metrics: 1) proportion of infected nodes, 2) average number of secondary infections, and 3) time-to-infection. We simulate the SIR model on networks to estimate the disease metrics in the samples and the UN. Here, we present the results for $\gamma=1$; see C.1 for $\gamma\in\{1/7,1/14\}$.
@@ 3.2.1 Proportion of infected nodes
Figure 4: The ratio of the proportion of infected nodes during an epidemic between RW (dashed red) and MHRW (dotted blue) samples, relative to their respective UNs, is shown for retained (1) and removed duplicate nodes (2). The light red and blue bands indicate the range of one standard deviation. The UN infection count is based on SIR model simulations ($\gamma=1$) for $10,000$ networks of each type (the parameters are found in Table 1) generated with the NetworkX library (Hagberg et al., 2008). Sample infection counts are derived from $100$ samples (size $500$) for $10,000$ networks using RW and MHRW.
@@
For $\beta>0.2$, RW samples consistently overestimate the proportion of infected nodes in ER and SW networks compared to MHRW samples, regardless of the inclusion of duplicate nodes (Figure 4). In SF networks, while MHRW samples tend to overestimate the average number of infected nodes relative to RW samples, estimates from both algorithms exhibit substantial variability and fail to represent the underlying network structure accurately.
For ER networks, the MHRW samples overestimate the proportion of infected nodes when duplicate nodes are removed (Figure 3(g)) but align well when duplicate nodes are retained (Figure 3(d)). RW estimates remain consistent regardless of duplicate removal. Similar behaviour is observed for SW networks in Figure 3(e) and Figure 3(h).
@@
We use hypothesis testing to statistically compare the estimates from the RW and MHRW sampling algorithms for the ER, SW, and SF networks, respectively. First, we conduct a one-sample KS test to assess whether the proportion of infected nodes follows a Normal distribution, as stated in Hypothesis 2.1, with significance level $\alpha=0.05$. The KS test statistic is summarized in B Table 6 for three data groups: UN, RW, and MHRW for ER, SW, and SF networks. The KS statistic is approximately $0.5$ across networks and data groups, with all corresponding p-values less than $0.05$. These results indicate significant deviation from normality, providing evidence to reject $H_{0}$ in Hypothesis 2.1.
Since the proportion of infected nodes does not follow a Normal distribution, we conduct a one-tailed U-test to compare the RW and the MHRW sample estimates with the UN. We test two hypotheses, one for statistically significant overestimation, Hypothesis 2.4, and another for statistically significant underestimation, Hypothesis 2.5, with significance level $\alpha=0.05$, for the ER, SW, and SF networks, respectively. The U-test statistic provides a measure of closeness between the sample estimates and the UN, with a relatively larger value indicating greater similarity to the UN. The statistics of the U-test are summarized in Table 2 for samples with duplicate nodes and samples without duplicate nodes.
The p-values are less than $0.05$ for three network types for the RW samples (with and without duplicates) and the MHRW samples without duplicate nodes, for Hypothesis 2.4. This provides evidence to reject $H_{0}$ in Hypothesis 2.4 and indicates the significant overestimation of the proportion of infected nodes. For the MHRW samples with duplicate nodes, the p-value is less than $0.05$ for the SF network, evidence to reject $H_{0}$ in Hypothesis 2.4. This indicates the significant overestimation by the MHRW samples for the SF network. On the other hand, for the MHRW samples with duplicate nodes in the ER and SW networks, the p-values are less than $0.05$ for Hypothesis 2.5 (reject $H_{0}$), indicating significant underestimation of the proportion of infected nodes.
@@
Table 2: U-test statistics (Hypotheses 2.4 and 2.5) for comparing the estimates of the proportion of infected nodes in the samples generated using the RW and the MHRW sampling algorithms with the UN, with and without duplicate nodes.
@@ 3.2.2 Average secondary infections
Figure 5: The ratio of the number of secondary infections during an epidemic between RW (dashed red) and MHRW (dotted blue) samples, relative to their respective UNs, is shown for (1) retained and (2) removed duplicate nodes. The light red and blue bands indicate the range of one standard deviation. The UN estimates are based on SIR model simulations ($\gamma=1$) for $10,000$ networks of each type (the parameters are found in Table 1) generated with the NetworkX library (Hagberg et al., 2008). Sample estimates are derived from $100$ samples (size $500$) for $10,000$ networks using RW and MHRW.
@@
For $\beta>0.3$, RW samples (with or without duplicates) overestimate the number of secondary infections in ER and SW networks compared to the MHRW samples as shown in Figures 4(a), 4(b), 4(d), and 4(e).
@@
For ER networks, the removal of duplicate nodes results in overestimation by MHRW samples (Figure 4(d)), while it aligns well when duplicate nodes are retained (Figure 4(a)). However, no change is observed for RW estimates in both scenarios (with or without duplicate nodes). Similar behaviour is observed for SW networks in Figure 4(b) and Figure 4(e).
@@
For the SF network, both sampling algorithms overestimate the number of secondary infections with high variability (Figures 4(c) and 4(f)). However, when duplicate nodes are retained, MHRW sample estimates are closer to the UN than the RW estimates, whereas this behaviour reverses when duplicates are removed. We consistently observe that the network properties and disease metrics are highly variable and more sensitive to changes in the system parameters for the SF networks, indicating the limitation of both sampling algorithms to represent the UN well.
@@
We conduct hypothesis testing to statistically compare the estimates from the sampling algorithms. First, we conduct the one-sample KS test to assess whether the number of secondary infections follows a Normal distribution, as stated in Hypothesis 2.2, with significance level $\alpha=0.05$, for the ER, SW, and SF networks, respectively. The KS test statistic is summarized in B Table 6 for three data groups that are UN, RW, and MHRW for ER, SW, and SF networks. The KS statistic ranges from $0.5$ to approximately $0.6$ across networks and data groups, with all corresponding p-values less than $0.05$. These results indicate significant deviation from normality, providing evidence to reject $H_{0}$ in Hypothesis 2.2.
@@
Since the number of secondary infections does not follow a Normal distribution, we conduct a one-tailed U-test to compare the RW and the MHRW sample estimates with the UN, for the ER, SW, and SF networks, respectively. We test two hypotheses, one for statistically significant overestimation (Hypothesis 2.6) and another for statistically significant underestimation (Hypothesis 2.7), with significance level $\alpha=0.05$. The statistics of the U-test are summarized in Table 3 for samples with duplicate nodes and samples without duplicate nodes.
@@
The p-values are less than $0.05$ individually for three network types for the RW samples (with and without duplicates) and the MHRW samples without duplicate nodes, for Hypothesis 2.6. This provides evidence to reject $H_{0}$ in Hypothesis 2.6 and indicates a significant overestimation of the number of secondary infections. For the MHRW samples with duplicate nodes, p-values are less than $0.05$, respectively for the ER and SF networks, evidence to reject $H_{0}$ in Hypothesis 2.7, implying a significant underestimation of the number of secondary infections. However, for the SW network, the p-value is greater than $0.05$ (accept $H_{0}$) for the MHRW sample estimates with duplicate nodes, for Hypotheses 2.7 and 2.6, indicating that the MHRW estimates are statistically similar to the UN.
@@
Similar to the proportion of infected nodes, in the case of the number of secondary infections, the U-test statistics (see Table 3) are higher for the MHRW sample estimates (with or without duplicates) compared to the RW for the ER and SW networks, implying that MHRW estimates are closer to the UN. For the SF networks, the RW sample estimates are closer to the UN compared to the MHRW, indicated by the higher values of U-test statistics in Table 3.